OpenStudy (agent_a):

Volume of Cylindrical Shells? Use the method of cylindrical shells to find the volume generated by rotating the region bounded by the given curves about the specified axis. y = x^4 , y = 0 , x = 1 ; about x=2 I've tried various formulae, but none of them seem to yield the correct answer. I know the correct answer, just not the solution. So far, I am veering towards V = ∫2π(x-2)(x^4)dx [integrated from 1 to 2], but the final answer does not match with the answer given by the book.

OpenStudy (isaiah.feynman):

Were you given the limits of integration or you solved for it?

OpenStudy (agent_a):

I solved for the limits of integration.

OpenStudy (isaiah.feynman):

And you're sure its correct?

OpenStudy (agent_a):

Nope. I'm only sure about the limits of integration, if the rotation was about the x-axis or y-axis. I know you have to account for the radius on how it is not solely x or y respectively, but rather, you have to subtract a certain number from the radius, because the rotation is done on a line. From what I understand, it should be x-2, because you rotate on x =2, but that doesn't seem to be the case.

OpenStudy (isaiah.feynman):

Its not x-2 I think the radius is 2.

OpenStudy (agent_a):

But if 2 is the radius, the cylinder would have to extend towards the y-axis, until one of its sides is flush with it. Initially, I was integrating from 0 to 2, but I realized that that would result in the radius' switch, completely. Because, under those circumstances, 2 would not be the radius anymore, and 1 would be the axis of rotation.

OpenStudy (agent_a):

Either that, or it is a larger cylinder (in the instance that 2 is the radius), but that cannot happen because the region of x =1 and x = 2 will be surpassed... not to mention, there are no points to intersect with y=x^4 thereafter, because y=x^4 only intersects x =2 at (2, 16).

OpenStudy (agent_a):

I know that you have to subtract 2 from x, because the cylinder does not cover the area from 0 to 1... That's how you account for the difference... at least, that's what I've picked-up, from scanning through various material.

OpenStudy (isaiah.feynman):

Tricky

OpenStudy (agent_a):

Thanks for trying, though.

OpenStudy (isaiah.feynman):

No problems.

OpenStudy (isaiah.feynman):

But I haven't stopped trying, when I get it I will tell you. Wait what is the correct answer to this problem?

OpenStudy (agent_a):

\[\frac{ 7 }{ 15 } \pi \]

OpenStudy (isaiah.feynman):

Ok thanks.

zepdrix (zepdrix):